# A Cognitive Evaluation Benchmark of Image Reasoning and Description for Large Vision-Language Models

### v3

@@ Abstract
Large Vision-Language Models (LVLMs), despite their recent success, are hardly comprehensively tested for their cognitive abilities. Inspired by the prevalent use of the “Cookie Theft” task in human cognition test, we propose a novel evaluation benchmark to evaluate high-level cognitive ability of LVLMs using images with rich semantics 111 The code and the data of CogBench is released at: https://github.com/xiujiesong/Cog-Bench. . It defines eight reasoning capabilities and consists of an image description task and a visual question answering task. Our evaluation on well-known LVLMs shows that there is still a large gap in cognitive ability between LVLMs and humans.
@@ 1 Introduction
Recently, with the emergence of Large Language Models (LLMs) such as GPT-4 [1], the cognitive abilities of language models have reached a new level [2]. They demonstrated remarkable performance in many tasks [3]. In Vision Language (VL), several researchers [4, 5, 6] endeavor to boost Vision-Language Pre-trained Models (VLPMs) by integrating powerful LLMs [7, 8], referred to as Large Vision-Language Models (LVLMs). With LLM as the “brain”, the cognitive abilities of LVLMs are also improved and more challenging tasks, such as table, chart, and document reasoning, can be solved [9]. Some state-of-the-art LVLMs, such as GPT-4V and GPT-4o[1], are progressing towards human-level cognitive abilities [9]. Thus, there’s a growing interest in evaluating the cognitive abilities of LVLMs. Such evaluation usually amounts to co-reasoning of images and text, e.g., question answering about a picture. Though some LVLM evaluation benchmarks, such as MME [10], MMBench [11], and SEED Bench [12], evaluate cognitive reasoning ability as one aspect of their evaluation, they do not provide a comprehensive evaluation of higher-level reasoning abilities and most of the images they use require relatively little reasoning to understand.
In our study, we draw inspiration from the “Cookie Theft” picture description task (Figure 3), a key component of the Boston Diagnostic Aphasia Examination [13], widely utilized in clinical practice within speech-language pathology [14] and cognitive function screening [15]. Notably, despite being designed over half a century ago, this picture remains prevalent in contemporary psychological discussions.
@@
Figure 1: Cookie Theft picture description task. The descriptions in the green frame and the orange frame were respectively produced by a 75-year-old healthy man and a 66-year-old woman with probable Alzheimer’s dementia [^3] The two samples are extracted from DementiaBank (https://dementia.talkbank.org/), which records description transcripts from both healthy subjects and dementia patients..
Figure 2: The comparison between our images and those from the previous visual reasoning tasks. Compared to our images, image (a) has fewer entities and CoRs, image (b) and (c) have some entities, but fewer CoRs.
@@
Is it possible to transfer the success of Cookie Theft picture description in human cognition test to evaluating cognitive abilities of LVLMs? Linguists and psychologists [14] conducted an analysis to determine the factors contributing to the success of “Cookie Theft”. The study revealed that the narrative includes information with varying levels of importance and encompasses a broad range of semantic categories. It was observed that during the description of “Cookie Theft”, individuals with intact cognitive functions exhibit their cognitive prowess by logically deducing the events and their interconnections. In contrast, those with cognitive impairments tend to merely list the superficial aspects of the situation without deeper reasoning. For instance, in Figure 3, by comparing the descriptions produced by the healthy man and the woman with probable Alzheimer’s dementia, we can identify the following differences:
@@
  • The description produced by a healthy man uses “mother” instead of “lady”, indicating the reasoning about character relationship.
@@
  • The healthy man also used “the mother is preoccupied” and “happy” to describe people’s mental state.
  • The description also reflects the causal relationships between events. Because “the mother is preoccupied” and “not paying attention to her kids”, “kids are stealing cookies.
Through these reasoning processes, the difference of cognitive abilities between the two individuals is reflected in their descriptions. A picture that can evaluate cognitive functions needs to be carefully designed and crafted. Tasnim et al. [16] introduced guidelines for drawing pictures similar to Cookie Theft, which is consistent with findings mentioned above. Generally speaking, compared to ordinary images, Cookie Theft-like images feature i) a prominent story theme, ii) richer content, iii) display complex relationships among entities, and thus require stronger cognitive abilities to understand and describe.
With the above design principles, we propose to construct a Cognitive Evaluation Benchmark, named as CogBench, which consists of high quality Cookie Theft-like images to evaluate the cognitive reasoning ability of LVLMs. CogBench defines eight core cognitive reasoning capabilities, including reasoning about special time, location, character, character relationship, event, event relationship, next moment event and mental state. Both a generative Image Description task and a discriminative Visual Question Answering (VQA) task are designed. Our main contributions are as follows:
  • To the best of our knowledge, this is the first-of-its-kind attempt to incorporate a cognitive test designed for humans into the evaluation of LVLMs.
  • We created and open-sourced the largest VL dataset with Cookie Theft-like images for evaluating LVLM’s cognitive abilities.
  • Our evaluation on existing LVLMs shows that a large gap exists between the cognitive ability of LVLMs and human beings, indicating CogBench will be a valuable evaluation benchmark in the near future.
@@ 2 Dataset Construction
In this section, we will introduce the construction of CogBench. We will introduce image collection, annotation, tasks and the statistics of CogBench. Ethical considerations of CogBench are discussed in Appendix G.
@@ 2.1 Image Collection
Based on previous studies [14, 16], we set the following image collection criteria:
  • Rule 1: Story-telling The image depicts an interesting story. For instance, the Cookie Theft picture tells the story of a mother busy washing dishes while two kids takes the opportunity to stand on a stool and sneakily steal cookies.
  • Rule 2: Rich Chain-of-Reasonings Images should display rich Chain-of-Reasonings (CoRs) in a scene. A CoR connects low-level observations in an image to produce a high-level reasoning conclusion or connects the cause and effect of events. For example, “The mother is busy washing dishes. + The boy is standing on the stool behind the mother. + The girl standing by the boy is shushing him. + The boy is fetching cookies from the jar in the cabinet. $\rightarrow$ The boy and girl are stealing cookies.” is a CoR about high-level event “stealing cookies”. Note that the story is actually constructed via these CoRs.
  • Rule 3: Content Simplicity Images should contain rich content but not be overly complex. The number of subjects should be limited for better emphasis on the key points of the story.
With the above criteria, we manually collect pictures from Pinterest [^4] https://www.pinterest.com/, and the Cookie Theft is also included in CogBench. Figure 2 shows the differences between our images and those from other datasets by examples.
@@ 2.2 Image Annotation
Figure 3: An example of Description task from CogBench.
@@
Human annotators, mostly undergraduate or graduate students aged 18-28, are hired to annotate the collected images [^5] Wages and recruitment process details can be found in Appendix G.. As shown in Figure 3, the annotation includes three parts: [Entities], [CoRs] and [Description]. By annotating [Entities] and [CoRs], we aim to evaluate the low-level recognition ability and high-level cognitive reasoning ability of models respectively based on description. [Description] is annotated as the reference description for the image. The three parts are annotated in that order.
@@
Entity Annotation We ask annotators to list as many entities in the image as possible and entities that are difficult to recognize should be omitted.
CoR Annotation In order to evaluate model cognition in a fine-grained manner, the following eight reasoning capabilities are annotated with CoRs:
  • Special Time Reasoning: reasoning about the special time of the story in the image, e.g., festivals, seasons etc.
  • Location Reasoning: reasoning about the location of the story in the image, e.g., near the school etc.
  • Character Reasoning: reasoning about the character of the subjects in the image, e.g., police officer etc.
@@
  • Character Relationship Reasoning: reasoning about relationship between characters in the image, e.g., “the woman is the mother of the kids.”
@@
  • Event Reasoning: reasoning about high-level events in the current moment and previous moments in the image based on clues in the picture. The difference between high-level and low-level lies in how much semantic information the event contains, e.g., “stealing cookies” is a higher-level event compared to “taking cookies” as it additionally conveys the semantic of “taking advantage without permission or knowledge.”
@@
  • Event Relationship Reasoning: reasoning about causal and temporal relationship between different events in the image. For instance, “the sink is overflowing because the mother left the tap on.”
@@
  • Next Moment Event Reasoning: reasoning about the event that will happen in the next moment. For example, “The police officer will reprimand the boy who violates the rules.”
  • Mental State Reasoning: reasoning about mental state of subjects in the image, e.g., daydreaming etc.
@@
Description Summary Annotators are last asked to write a description to tell the entire story in the image based on the annotated [Entities] and [CoRs].
@@
The complete annotation instruction is in Appendix A. Considering different people may have different understanding about some images, we ask three annotators to annotate each image. Then, we manually merge the three annotations as one final annotation by majority vote. For the [Description], if two or more annotators understand the image in the same way, we accept the best descriptions as the final description. For [Entities] and [CoRs], we first accept most of the entities and CoRs that are annotated by at least two annotators. Other entities and CoRs are also included if they are reasonable. For images with significantly different understandings from the three annotators, we will discard these images.
@@ 2.3.1 Image Description Task
This is the primary task of the benchmark. The difference between our description task and existing image description task [17, 18, 19] is that we expect LVLMs understand and describe the story in the image through high-level cognitive reasoning. For instance, in Figure 3, the description of the image should not only include what is in the picture but also focus on elucidating the story of “parents are going out to attend a formal event, and the mother is refusing the boy to accompany them” through a series of reasonings.
@@ 2.3.2 Visual Question Answering Task
Figure 4: Generating a MCQ from event reasoning annotations.
The VQA task features standard four-option Multiple-Choice Questions, easing the evaluation process. Like the Description task, VQA questions involve different types of high-level cognitive reasoning, as illustrated by the question about events in Figure 4.
@@ 2.4 CogBench Statistics
CoR 47 179 106 263 701 425 107 417
QA 86 220 162 317 658 402 135 597
CogBench consists of 251 semantically-rich images with a total of 2670 entities, 2245 CoRs, 251 descriptions and 2577 questions, indicating the content contained in each image is complex, showcased in Table 1. The number of CoRs in event-related reasoning and [Mental State Reasoning] is large, which is a manifestation of the rich interesting stories in the images.
@@ 3 Experiments
In this section, we will evaluate selected LVLMs, detail evaluation metrics and discuss the results.
@@ 3.1 Large Vision-Language Models
We evaluated a selection of recent representative open-source and closed-source LVLMs in this paper, including InstructBLIP-7B [20], Qwen-VL-Chat [21], LLaVA-v1.5-7B and LLaVA-v1.5-13B [22], mPLUG-Owl-2 [6], ShareGPT4V-7B and ShareGPT4V-13B [23], GPT-4V and GPT-4o [1]. A brief introduction of these models is in Appendix D. These open-source models are deployed on a machine with four GeForce RTX 2080 Ti GPUs and a machine with one GeForce RTX 3090 GPU.
@@ 3.2 CogBench Evaluation Strategy
Evaluation for Description Task
Here we consider model performance at two levels: low-level Recognition ability and high-level Cognition ability. Evaluation metrics for both levels are calculated using recall scores, referred to as Recognition Score and Cognition Score, respectively.
The Recognition Score is calculated as the ratio of the number of recognized [Entities] to the number of annotated [Entities] in all images. First, we use SpaCy [^6] https://spacy.io/ to extract nouns from the model generated description, and then calculate cosine similarity between embeddings [^7] Implemented with sentence-transformers package (https://www.sbert.net) and all-mpnet-base-v2 is adopted as the model to encode [Entities] and nouns. of annotated [Entities] and extracted nouns. For each entity, if the cosine similarity score between the entity and any noun is greater than a threshold (0.6 in this paper), we consider the entity is recognized by the model.
For Cognition Score, we calculate both scores for each of the eight cognitive reasoning abilities and an overall score using GPT-4 [1]. To enhance objectivity and granularity, GPT-4 is utilized for a binary classification task to assess if generated descriptions include the semantics of annotated CoRs, as detailed in Appendix C. The CoR scores for each type of reasoning capability are then used to compute a recall score for each respective type. The overall Cognition Score is the sum of all CoR scores divided by the total number of CoRs.
Evaluation for VQA Task
For Multiple-Choice Questions in VQA task, we use accuracy as the evaluation metric. As questions are generated based on CoRs, we can both calculate the accuracy for each reasoning capability and the overall cognition capability.
3.3 Results of Description Task
We prompt the selected LVLMs with the following instruction to obtain detailed descriptions about images in CogBench: ‘‘Describe this image in detail.’’ We evaluate the performance of LVLMs on Description task in terms of both recognition and cognition ability. As a reference, we also calculated traditional image captioning evaluation metrics by comparing annotated reference description and model-generated description, and details are shown in Appendix E.

### v4

@@ Abstract
Large Vision-Language Models (LVLMs), despite their recent success, are hardly comprehensively tested for their cognitive abilities. Inspired by the prevalent use of the Cookie Theft task in human cognitive tests, we propose a novel evaluation benchmark to evaluate high-level cognitive abilities of LVLMs using images with rich semantics. The benchmark consists of 251 images along with comprehensive annotations. It defines eight reasoning capabilities and comprises an image description task and a visual question answering task. Our evaluation of well-known LVLMs shows that there is still a significant gap in cognitive abilities between LVLMs and humans111 Our code and data are available at: https://github.com/xiujiesong/CogBench. .
A Cognitive Evaluation Benchmark of Image Reasoning and Description for Large Vision-Language Models
Xiujie Song1, Mengyue Wu1thanks: Corresponding authors., Kenny Q. Zhu211footnotemark: 1, Chunhao Zhang1, Yanyi Chen3 1 X-LANCE Lab, Department of Computer Science and Engineering MoE Key Lab of Artificial Intelligence, AI Institute Shanghai Jiao Tong University, Shanghai, China 2University of Texas at Arlington, Arlington, Texas, USA 3University of Chicago, Chicago, Illinois, USA 1{xiujiesong, mengyuewu}@sjtu.edu.cn, 2kenny.zhu@uta.edu
@@ 1 Introduction
Recently, with the emergence of Large Language Models (LLMs) such as GPT-4 OpenAI (2023), the cognitive abilities of language models have reached a new level Zhuang et al. (2023). They demonstrate remarkable performance in many tasks Bubeck et al. (2023). In Vision-Language (VL), several researchers Zhu et al. (2023b); Liu et al. (2023b); Ye et al. (2023) endeavor to boost Vision-Language Pre-trained Models (VLPMs) by integrating powerful LLMs Touvron et al. (2023); Chiang et al. (2023), referred to as Large Vision-Language Models (LVLMs) Li et al. (2023c). With LLM serving as the “brain”, the cognitive abilities of LVLMs are enhanced, enabling them to tackle more challenging tasks Yang et al. (2023). Some state-of-the-art LVLMs, such as GPT-4o OpenAI (2023), are progressing toward human-level cognitive abilities. Thus, there is growing interest in evaluating cognitive abilities of LVLMs. Though some LVLM evaluation benchmarks, such as MME Fu et al. (2023), MMBench Liu et al. (2023c), and SEED Bench Li et al. (2023a), evaluate cognitive reasoning ability as one aspect of their evaluation, they do not provide a comprehensive evaluation of higher-level reasoning abilities. Most of the images they use contain less semantics and thus require relatively little reasoning to understand.
@@
Figure 1: Cookie Theft picture description task. The descriptions in the green frame and the orange frame were respectively produced by a healthy 75-year-old man and a 66-year-old woman with probable AD dementia [^3] The two samples are extracted from DementiaBank (https://dementia.talkbank.org/), which records description transcripts from both healthy subjects and dementia patients..
In our study, we draw inspiration from the Cookie Theft picture description task (Figure 3), a key component of the Boston Diagnostic Aphasia Examination Goodglass et al. (2001), which is widely used in clinical practice within speech-language pathology for language and cognitive function screening Cummings (2019); Mueller et al. (2018). Notably, despite being designed more than half a century ago, this picture remains prevalent in contemporary psychological discussions.
@@
Is it possible to transfer the success of the Cookie Theft picture description in human cognitive tests to evaluating the cognitive abilities of LVLMs? Linguists and psychologists (Cummings, 2019) conducted an analysis to determine the factors contributing to the success of the Cookie Theft picture. The study reveals that the narrative includes information with varying levels of importance and encompasses a broad range of semantic categories. It is observed that during the description of the Cookie Theft picture, individuals with intact cognitive functions exhibit their cognitive prowess by logically deducing the events and their interconnections, the relationships between characters and their mental states, etc. In contrast, those with cognitive impairments tend to merely list the superficial aspects of the situation without deeper reasoning. For instance, in Figure 3, by comparing the descriptions produced by the healthy man and the woman with probable Alzheimer’s disease (AD) dementia, we can identify the following differences:
@@
  • The description produced by the healthy man used “mother” instead of “lady”, indicating reasoning about character relationship.
@@
  • The healthy man used “the mother is preoccupied” and “happy” to describe people’s mental state.
  • The description reflected the causal relationships between events. “The kids are stealing cookies”, because “the mother is preoccupied” and “not paying attention to her kids”.
Through these reasoning processes, the difference in cognitive abilities between the two individuals is reflected in their descriptions. A picture that can evaluate cognitive functions needs to be carefully designed and crafted. Tasnim et al. (2022) introduced guidelines for drawing pictures similar to Cookie Theft, which is consistent with the findings mentioned above. Generally speaking, compared to ordinary images, Cookie Theft-like images feature i) a prominent story theme, ii) richer content, iii) display complex relationships among entities, and thus require stronger cognitive abilities to understand and describe.
With the above design principles, we propose to construct a Cognitive Evaluation Benchmark, referred to as CogBench, to evaluate cognitive abilities of LVLMs mainly from the reasoning perspective using high-quality Cookie Theft-like images. CogBench defines eight core cognitive reasoning capabilities, including reasoning about special time, location, character, character relationship, event, event relationship, next moment event and mental state. Both a generative Image Description task and a discriminative Visual Question Answering (VQA) task are designed. Our main contributions are as follows:
  • To the best of our knowledge, this is the first-of-its-kind attempt to incorporate the concept of the well-known Cookie Theft picture description task, originally designed for human cognitive testing, into the cognitive evaluation of LVLMs.
  • Taking insights from human cognition research, we are the first to define Cookie Theft-like images with eight reasoning dimensions and to create a dataset with semantically complex images aligned with the Cookie Theft. This approach allows for a comprehensive evaluation of the visual reasoning capabilities of LVLMs across these dimensions.
  • Our evaluation on existing LVLMs shows that a significant gap exists between the cognitive abilities of LVLMs and human beings, indicating CogBench will be a valuable evaluation benchmark in the near future.
@@ 2 Dataset Construction
In this section, we will introduce the construction of CogBench, detailing its image collection, annotation process, tasks, and data statistics.
@@ 2.1 Image Collection
Figure 2: Comparison between our images and those from previous visual reasoning tasks. Our images contain rich entities and CoRs. Compared to our images, image (a) has fewer entities and CoRs, while image (b) and (c) have some entities but fewer CoRs.
Building on previous studies Cummings (2019); Tasnim et al. (2022), we establish the following criteria for collecting Cookie Theft-like images that we propose:
Rule 1: Storytelling
The image depicts an interesting story. For instance, the Cookie Theft picture tells the story of a mother busy washing dishes while two kids take the opportunity to stand on a stool and sneakily steal cookies.
Rule 2: Rich Chain-of-Reasonings
Images should display rich Chain-of-Reasonings (CoRs) in a scene. A CoR connects low-level observations in an image to produce a high-level reasoning conclusion or connects the cause and effect of events. For example, “The mother is busy washing dishes. + The boy is standing on the stool behind the mother. + The girl standing by the boy is shushing him. + The boy is fetching cookies from the jar in the cabinet. $\rightarrow$ The boy and girl are stealing cookies.” is a CoR about the high-level event “stealing cookies”. Note that a story is usually constructed through various CoRs.
Rule 3: Restricted Content Complexity
Images should contain rich content but not be overly complex. The number of entities should be sufficient to support a good story while being restricted to emphasize the key points effectively.
With the above criteria, we manually collect pictures from Pinterest [^4] https://www.pinterest.com/, and the Cookie Theft picture is also included. Figure 2 shows the differences between our images and those from other datasets by examples. Note that most of the images in CogBench are in a painting style because they are abstracted from reality and tend to contain richer CoRs than real-world images, making them more effective for picture description tasks. Even so, painting-style images that meet our criteria are still rare, highlighting the value of our dataset.
@@ 2.2 Image Annotation
Figure 3: An example of the Description task from CogBench.
@@
Human annotators, mostly undergraduate or graduate students aged 18-28, are hired to annotate the collected images. As shown in Figure 3, the annotation includes three parts: [Entities], [CoRs] and [Description]. By annotating [Entities] and [CoRs], we aim to evaluate the low-level recognition ability and high-level cognitive reasoning ability of models respectively based on their descriptions. [Description] is annotated as the reference description for the image. The three parts are annotated in that order.
@@
[[[[Entity Annotation]]]] We ask annotators to list as many entities in the image as possible and entities that are difficult to recognize should be omitted.
[[[[CoR Annotation]]]] To evaluate model cognition in a fine-grained manner, we design eight reasoning dimensions based on studies in human cognition Cummings (2019); Byom and Mutlu (2013); Addis et al. (2007) and Computer Vision (CV) Fu et al. (2022); Park et al. (2020); Zellers et al. (2019); Shen et al. (2020). CoRs for these dimensions are annotated:
  • Special Time Reasoning: reasoning about the special time of the story in the image, e.g., festivals.
  • Location Reasoning: reasoning about the location of the story in the image, e.g., near a school.
  • Character Reasoning: reasoning about the characters of subjects in the image, e.g., a doctor.
@@
  • Character Relationship Reasoning: reasoning about the relationships between characters in the image, e.g., “the woman is the mother of the kids.”
@@
  • Event Reasoning: reasoning about the high-level events in the current and previous moments in the image. The difference between high-level and low-level events lies in the amount of semantic information they contain. For example, “stealing cookies” is a higher-level event compared to “taking cookies” as it additionally conveys the semantic of “taking advantage without permission or knowledge.”
@@
  • Event Relationship Reasoning: reasoning about the causal and temporal relationships between different events in the image. For instance, “the sink is overflowing because the mother left the tap on.”
@@
  • Next Moment Event Reasoning: reasoning about the events that will happen in the next moment in the image. For example, “the police officer will reprimand the boy who violates the rules.”
  • Mental State Reasoning: reasoning about the mental states of subjects in the image, including their emotions, thoughts, and other psychological states. For example, “the girl appears to be happy.”
@@
[[[[Description Summary]]]] Annotators are finally asked to write a description that conveys the entire story in the image based on the annotated [Entities] and [CoRs].
@@
The complete annotation instruction can be found in Appendix A. Considering different people may have different understanding about some images, we ask three annotators to annotate each image. Then, we draw on the idea of majority voting to merge the three annotations into one. For [Entities] and [CoRs], we first accept most of the entities and CoRs that are annotated by at least two annotators. Other entities and CoRs are also included if reasonable. The final [Description] is obtained by modifying the best annotated [Description] with the merged [Entities] and [CoRs]. We discard images where there is a significant difference in understanding among the three annotators.
@@ 2.3.1 Image Description Task
This is the primary task of the benchmark. The difference between our description task and existing image description tasks Xie et al. (2022); Zhu et al. (2023a); Zhuge et al. (2023) is that we expect LVLMs to understand and describe the story in the image through high-level cognitive reasoning. For instance, in Figure 3, the description of the image should not only include what is in the picture but also focus on elucidating the story of “on a cold winter day, a man is reading a newspaper in a phone booth near the bus stop to escape the cold, while two passing women express their displeasure upon seeing that” through a series of reasoning processes.
@@ 2.3.2 Visual Question Answering Task
Figure 4: Generating a multiple-choice question based on an [Event Reasoning] CoR annotation.
@@ 2.4 CogBench Statistics
CogBench consists of 251 semantically rich images with a total of 2670 entities, 2243 CoRs, 251 descriptions and 2577 questions, indicating the content contained in each image is complex, showcased in Table 1. The number of CoRs of event-related reasoning and [Mental State Reasoning] is large, which is a manifestation of the rich interesting stories in the images.
@@ 3 Experiments
We will evaluate the selected LVLMs, detail evaluation strategies and discuss the results in this section.
@@ 3.1 Large Vision-Language Models
We evaluate a selection of recent representative open-source and closed-source LVLMs, including InstructBLIP-7B Dai et al. (2023), Qwen-VL series models Bai et al. (2023); Wang et al. (2024), mPLUG-Owl-2 Ye et al. (2023), LLaVA series models Liu et al. (2023a, 2024); Li et al. (2024), ShareGPT4V Chen et al. (2023a), CogVLM series models Wang et al. (2023); Hong et al. (2024), InternVL2-26B Chen et al. (2023c), GPT-4V and GPT-4o OpenAI (2023). A brief introduction to these models is provided in Appendix C.
@@ 3.2 CogBench Evaluation Strategy
3.2.1 Evaluation of the Description Task
Evaluation Modes
For the Description task, we set up two evaluation modes: Spontaneous Description and Directed Reasoning. In the Spontaneous Description mode, we prompt the LVLMs with the following instruction to obtain detailed image descriptions: “Describe this image in detail.” This mode is more similar to the Cookie Theft picture description task, which aims to stimulate spontaneous descriptions Matias-Guiu et al. (2022). It can help analyze the behavior of LVLMs when they describe images. For the Directed Reasoning mode, the corresponding prompt is: “Please provide a detailed description of the story depicted in the image, including high-level reasoning about the time and location, the roles and relationships of the characters, the events and their causal relationships, what might happen next, and the mental states of the characters.” This mode simplifies the Description task compared to the Spontaneous Description mode and aims to evaluate whether models can reason correctly when they know the directions we expect.
Evaluation Metrics
We consider model performance at two levels: low-level Recognition ability and high-level Cognition ability. Evaluation metrics for both levels are calculated based on recall scores, referred to as Recognition Score and Cognition Score, respectively.
The Recognition Score is calculated as the ratio of recognized [Entities] to annotated [Entities] across all images. First, we use spaCy [^5] https://spacy.io/ to extract nouns from the model-generated description, and use sentence-transformers [^6] https://www.sbert.net. The model all-mpnet-base-v2 is adopted. to encode the annotated [Entities] and extracted nouns into embeddings. Then, we calculate the cosine similarity between the embeddings of the [Entities] and the nouns. For each entity, if the cosine similarity score between the entity and any noun is greater than a threshold (0.6 in this paper), we consider the entity to be recognized by the model.
For the Cognition Score, we calculate the scores for each of the eight cognitive reasoning dimensions, as well as an overall score using GPT-4. To enhance objectivity and granularity, GPT-4 is utilized for a binary classification task to assess if a generated description includes the semantics of each annotated CoR. For reasoning types other than [Event Relationship Reasoning], we task GPT-4 with determining whether the conclusion in each CoR is mentioned in the description. For [Event Relationship Reasoning], we task GPT-4 with determining whether each causal relationship between events (i.e., the entire CoR), as annotated, is present in the description. The CoR scores for each dimension are then used to compute a recall score for each respective type. The overall Cognition Score is the sum of all CoR scores divided by the total number of CoRs. The corresponding prompts are shown in Appendix D.
3.2.2 Evaluation of the VQA Task
For multiple-choice questions in the VQA task, we use accuracy as the evaluation metric. As questions are generated based on CoRs, we can also calculate the accuracy for each reasoning capability as well as the overall cognitive capability.
3.3 Results of the Description Task
We evaluate the LVLMs on the Description task in terms of both recognition and cognition abilities. As a reference, we also calculate traditional image captioning evaluation metrics by comparing the model-generated description with the annotated reference [Description], and details are shown in Appendix E.